OpenStudy (anonymous):

Please help, precalc test tomorrow! Question is find all solutions on interval 0 less than or equal to theta less than or equal to 2 pi for 2sintheta+sqrt3=0 I'm confused on how to get 2theta.....

OpenStudy (anonymous):

Okay. 2sinθ+1 = 0 2sinθ = -1 sinθ = -1/2 sin π/6 = 1/2, so using that as a reference angle in QIII and QIV, we get the solutions π + π/6 and 2π - π/6

OpenStudy (anonymous):

Hope this helped. Have a great day!

OpenStudy (anonymous):

its not 2 sin theta, it's sin(2theta)